For the equation 4x – 2y = 6 find the
(a) slope 4/2
(b) y – intercept (0, -3)

Answer by sarah_adam(201)

y = mx + c
4x – 2y = 6
2y = 4x - 6
y = (4/2)x - 3
slope m = 4/2